OpenStudy (anonymous):

3x-2y ^2 would my answer be 9x^2 + 4y^2 im thinking that would be my final answer if it is not how can i figure it out

OpenStudy (uri):

Is it (3x-2y)^2?

OpenStudy (anonymous):

yes

OpenStudy (uri):

(a-b)^2=a^2+b^2- 2ab

OpenStudy (uri):

Use this formula

OpenStudy (uri):

Then, so (3x-2y)^2=9x^2+4y^2-12xy

OpenStudy (uri):

That's the answer after *=* sign.

OpenStudy (anonymous):

thank u

OpenStudy (uri):

Understand @suzy4321

OpenStudy (anonymous):

i am mislead with the 2ab how can you add that or why

OpenStudy (uri):

a=3 b=2 So 2ab means 2a*b So 2(3)(2)=12

OpenStudy (uri):

2x3x2

OpenStudy (anonymous):

i didnt mean that but thanks

OpenStudy (uri):

Oh you mean why is 2ab there? Well because of the formula :P

OpenStudy (anonymous):

yeah i dont remember this formulas gotta go back ...thank u

OpenStudy (uri):

If the question is (3x-2y)^2 formula will be:(a-b)^2=a^2+b^2- 2ab If the question is (3x+2y)^2 formula will be:(a+b)^2=a^2+b^2+ 2ab

OpenStudy (anonymous):

better sense

OpenStudy (uri):

Depends on the sign.